OpenStudy (anonymous):

Which point lies in the solution set for the following system of inequalities? y < x + 4 y < -2x + 2 (2, 3) (0, -5) (0, 5) (1, 0)

OpenStudy (anonymous):

If we pluck in the points into the inequalities, we can see which one works if it makes both of the inequalities true. By doing this, only the 2nd answer works. -5 < 0 + 4 = -5 < 4 [This is true] Now we do the other problem- -5 < -2(0) + 2 = -5 < -2 + 2 = -5 < 0 [This is also true]

OpenStudy (anonymous):

ohh okay :) thank you